# A Systematic Review of NLP for Dementia- Tasks, Datasets and Opportunities

### v1

@@ 5 Conclusions
Our review of 205 papers on NLP for dementia, spanning multiple scientific communities, reveals a richness of tasks, datasets, gaps, and creative opportunities. We show that NLP for dementia extends far beyond detection, with the potential to transform the lives of caregivers and patients, while shaping medical research on the disease. We call on researchers to collaborate closely and share knowledge through rigorous research, so that together we can make a meaningful impact on people’s lives and combat this disease.
@@ A.1 Search Methodology
We searched in six data sources: ACL Anthology, PubMed, DBLP, IEEE Xplore, Springer, and Wiley. The first database search was conducted on April 10, 2024, and the last one on July 19, 2024. Our research strategy focused on papers addressing dementia, Alzheimer’s, and Mild Cognitive Impairment (MCI) using NLP, including but not limited to large language models (LLMs). We searched for papers including the keywords ("NLP" OR "Natural Language Processing" OR "LLM" OR "Language Model" OR "ChatGPT") AND ("Dementia" OR "Alzheimer" OR "MCI" OR "Cognitive Impairment") in their title and/or abstract. Although we don’t include keywords for specific dementia pathologies (e.g., ’Lewy Body’), we believe that using the general term ’dementia’ in the title and/or abstract will capture papers on various types.
@@ A.2 Paper inclusion and Exclusion Criteria
Works applying NLP technologies, algorithms, and methodologies to textual data (including speech transcriptions) were included.
@@
Multi-modal approaches, as long as they include some form of textual analysis, were included.
@@
Only fully accessible, English-language papers were included.
e)
Only peer-reviewed papers were considered.
A.3 Dataset inclusion and Exclusion Criteria
We screened the featured datasets against the following criteria:
a)
Only datasets mentioned in the reviewed works were included.
b)
Textual data must be fully or partially in English.
c)
Sufficient information about the dataset must be available online or through private requests from researchers or institutions.
d)
Data does not need to be publicly accessible to be included in our review.
@@ A.4 Paper Annotation
Appendix B Extended Mention of Works
@@
In Section 3, we cited a representative sample of dementia detection works we reviewed. Below, we cite the remainder of the papers according to internal categorization used for our internal analysis.
@@
Detection works using classic ML classifiers on a wide range of features, from straightforward N-grams and information content units extracted from picture descriptions to dementia patients’ dental records and motor signs: Prud’hommeaux and Roark (2011), Fraser et al. (2013), Jarrold et al. (2014), Prud’hommeaux and Roark (2015), Fraser et al. (2016a), Yancheva and Rudzicz (2016), Zhou et al. (2016), Weissenbacher et al. (2016), Mirheidari et al. (2016), SANTOS et al. (2017), Wankerl et al. (2017), Pou-Prom and Rudzicz (2018), Al-Harrasi et al. (2021), Eyre et al. (2020), Chen et al. (2020), Searle et al. (2020), Martinc and Pollak (2020), Hane et al. (2020), Nasreen et al. (2021b), Gónzalez Atienza et al. (2021), Soni et al. (2021), Penfold et al. (2022), Ablimit et al. (2022b), Amini et al. (2023), Taghibeyglou and Rudzicz (2023), Patel and Wu (2024).
@@
Works using classic neural models such as LSTMs and transformer-based models like BERT: Karlekar et al. (2018), Pompili et al. (2018), Pou-Prom and Rudzicz (2018), Di Palo and Parde (2019), Chen et al. (2019), Edwards et al. (2020), Syed et al. (2020), Koo et al. (2020), Sarawgi et al. (2020), Pompili et al. (2020), Cummins et al. (2020), Balagopalan et al. (2020), Yuan et al. (2020), Pappagari et al. (2020), Haulcy and Glass (2021), Balagopalan et al. (2021), Wang et al. (2021), Qiao et al. (2021), Campbell et al. (2021), Rohanian et al. (2021a), Syed et al. (2021), Pappagari et al. (2021), Pérez-Toro et al. (2021), Mirheidari et al. (2021), Liu et al. (2021), Zhu et al. (2021), Saltz et al. (2021), Khan et al. (2022), Deng et al. (2022), Horigome et al. (2022), Zheng et al. (2022), Liu et al. (2022b), Liu et al. (2022a), Wang et al. (2022), Matošević and Jović (2022), Amini et al. (2023), Abdelhalim et al. (2023), Bouazizi et al. (2023), Gkoumas et al. (2023), Cai et al. (2023), Wang et al. (2023b), Shakeri et al. (2024), Dong et al. (2024), Kim et al. (2024).
@@
Works classifying between more granular categories such as ’MCI’, ’Alzheimer’s’, and ’Probable Alzheimer’s’: Mao et al. (2023), Liu et al. (2023b), Amini et al. (2024), Laurentiev et al. (2024). Works focusing on text-based MMSE score regression for dementia detection: Yancheva et al. (2015), Pou-Prom and Rudzicz (2018), Rohanian et al. (2021b).
@@
Works presenting multi-modal approaches (with audio / video) as well as augmentation methods using synthetic, noisy, cross-domain, or semantically and structurally different datasets: Che et al. (2017), Gligic et al. (2020), Gupta et al. (2020), Amin-Nejad et al. (2020), Guo et al. (2021), Zhu et al. (2021), Zhu et al. (2022a), Ablimit et al. (2022a), Mirheidari et al. (2022), Hlédiková et al. (2022), Ilias and Askounis (2022b), Maclagan et al. (2023), Chen et al. (2023b), Farzana and Parde (2023), Cui et al. (2023), Chen et al. (2023a), Cai et al. (2023), Lin and Washington (2024), Fard et al. (2024). Works leveraging LLMs for data augmentation: Bullard et al. (2016), Liu et al. (2019), Chintagunta et al. (2021), Liu et al. (2023a), Cai et al. (2023), Duan et al. (2023), Koga et al. (2024), Latif and Kim (2024).
@@ Appendix C Full Dataset Table
Table 2 (presented in a separate page due to margin requirements of the TACL format) showcases the dementia-related datasets encountered throughout our review. Datasets such as the Reagan Library and IMDB, though creatively used in some works, are excluded as they are not dementia-specific. Private health records or unreleased clinical notes are excluded as well.
@@
| Farmington Heart Study | x | x | - | - | - | - | x | x | - | x | x |
- | - | x | - | - | - | x | x | x | x | x | - |

### v2

@@ 8 Conclusions
Through a cohort of 242 papers, we uncovered the depth and richness of NLP research for dementia. Multiple scientific communities are working to advance early detection, support patients and caregivers, and much more- all through language. Yet, many gaps and opportunities remain, and the data is out there. We urge researchers to collaborate, share knowledge through rigorous research, and uphold the highest ethical standards to make a true impact and help combat this disease.
@@ A.1 Search Methodology
We searched in six data sources: ACL Anthology, PubMed, DBLP, IEEE Xplore, Springer, and Wiley. The database search was conducted between April and July 2024, with an additional search in January 2025 as part of our revised submission. Our research strategy focused on papers addressing dementia, Alzheimer’s, and Mild Cognitive Impairment (MCI) using NLP, including but not limited to large language models (LLMs). We searched for papers including the keywords ("NLP" OR "Natural Language Processing" OR "LLM" OR "Language Model" OR "ChatGPT") AND ("Dementia" OR "Alzheimer" OR "MCI" OR "Cognitive Impairment") in their title, abstract, and/or keywords. While we do not include keywords for specific dementia pathologies (e.g., Lewy body), we believe that using the general term dementia in the title, abstract and/or keywords will effectively capture papers on various types of the condition.
@@ A.2 Paper inclusion and Exclusion Criteria
Only fully accessible, English-language papers were included.
@@
Only peer-reviewed papers were included.
@@
Only full academic papers (excluding posters and theses) were included.
d)
Studies focusing on text as the primary modality or employing multi-modal approaches that incorporate textual analysis were included.
e)
@@
f)
Papers that, despite mentioning our defined keywords, do not directly deal with NLP applications to Dementia, were excluded.
Papers were deemed irrelevant to our literature review if they contained dementia- and NLP-related terms in their title, abstract, or keywords but focused on entirely unrelated topics. For example, Botros et al. (2020) mentions ‘dementia’ in the abstract and ‘language model’ in the keywords, but primarily focuses on sensor locations. Similarly, Daudet et al. (2016) includes ‘Alzheimer’s’ and ‘Language Model’ as keywords but primarily discusses the design of an application for diagnosing brain injuries using youth speech data.
As mentioned in Section 2, the first three steps of our screening process were automated, while the remaining steps - including gauging general relevance - were conducted manually by two PhD students from our NLP lab, one of whom has a medical background in Alzheimer-related projects. We encountered discrepancies in fewer than 10% of reviewed works, and resolved them through discussion until a final consensus was reached.
A PRISMA flowchart describing our search process can be seen in Figure 6.
A.3 Dataset inclusion and Exclusion Criteria
Throughout our review, we discuss classic and contemporary datasets, summarized in Table C. The datasets mentioned in the paper and listed in this table were screened against the following criteria:
a)
Only datasets referenced in the reviewed papers were included.
b)
The dataset must contain data that is fully or partially in English.
c)
Sufficient information about the dataset must be available online or obtainable through private requests to researchers or institutions.
@@ A.3 Dataset inclusion and Exclusion Criteria
Datasets do not need to be publicly accessible to be included in our review.
@@ A.4 Paper Annotation
Appendix B Extended Mention of Studies
@@
In Section 3, we cited a representative sample of dementia detection studies we reviewed. Below, we cite the remainder of the papers according to internal categorization used for our internal analysis.
@@
Detection papers using classic ML classifiers on a wide range of features, from straightforward N-grams and part-of-speech tagging to information content units extracted from picture descriptions and even dementia patients’ dental records and motor signs: Prud’hommeaux and Roark (2011); Fraser et al. (2013); Jarrold et al. (2014); Prud’hommeaux and Roark (2015); Fraser et al. (2016a); Yancheva and Rudzicz (2016); Zhou et al. (2016); Weissenbacher et al. (2016); Mirheidari et al. (2016); SANTOS et al. (2017); Wankerl et al. (2017); Klumpp et al. (2018); Pou-Prom and Rudzicz (2018); Eyre et al. (2020); Chen et al. (2020); Searle et al. (2020); Martinc and Pollak (2020); Hane et al. (2020); Al-Harrasi et al. (2021); Nasreen et al. (2021b); Clarke et al. (2021); Gónzalez Atienza et al. (2021); Soni et al. (2021); Penfold et al. (2022); Ablimit et al. (2022b); Dey and Mittal (2022); Amini et al. (2023); Liu and Wang (2023); Taghibeyglou and Rudzicz (2023); Wen et al. (2023); Patel and Wu (2024); Sharma et al. (2024)
@@
Studies using classic neural models such as LSTMs and transformer-based models like BERT: Karlekar et al. (2018), Pompili et al. (2018), Pou-Prom and Rudzicz (2018), Hong et al. (2019), Di Palo and Parde (2019), Chen et al. (2019), Fritsch et al. (2019), Pan et al. (2019), Edwards et al. (2020), Syed et al. (2020), Koo et al. (2020), Sarawgi et al. (2020), Pompili et al. (2020b), Cummins et al. (2020), Balagopalan et al. (2020), Yuan et al. (2020), Pappagari et al. (2020), Haulcy and Glass (2021), Valsaraj et al. (2021), Balagopalan et al. (2021), Wang et al. (2021), Qiao et al. (2021), Campbell et al. (2021), Nambiar et al. (2022), Khan et al. (2022b), Rohanian et al. (2021a), Syed et al. (2021), Pappagari et al. (2021), Pérez-Toro et al. (2021), Mirheidari et al. (2021), Liu et al. (2021), Zhu et al. (2021), Saltz et al. (2021), Khan et al. (2022a), Deng et al. (2022), Bouazizi et al. (2022), Horigome et al. (2022), Zheng et al. (2022), Liu et al. (2022b), Liu et al. (2022a), Wang et al. (2022), Matošević and Jović (2022), Amini et al. (2023), Rauniyar et al. (2023) Abdelhalim et al. (2023), Bouazizi et al. (2023), Gkoumas et al. (2023), Cai et al. (2023), Wang et al. (2023b), Shakeri et al. (2024), Dong et al. (2024), Kim et al. (2024), Laurentiev et al. (2024).
@@
Studies classifying between more granular stages of the dementia (e.g.‘MCI’, ‘Alzheimer’s’), assess disease progression or differentiate between dementia types (e.g., Alzheimer’s versus Frontotemporal), through classification: Thomas et al. (2005); Orimaye et al. (2016); Wang et al. (2020b); Padhee et al. (2020); Merone et al. (2022); Mao et al. (2023); Liu et al. (2023b); Amini et al. (2024); Laurentiev et al. (2024); Panahi et al. (2024). Studies focusing on text-based MMSE score regression for dementia detection: Yancheva et al. (2015), Pou-Prom and Rudzicz (2018), Farzana and Parde (2020), Rohanian et al. (2021b), Stoppa et al. (2023),Aryal et al. (2023).
@@
Studies presenting multi-modal approaches (with audio / video) as well as augmentation methods using synthetic, noisy, cross-domain, or semantically and structurally different datasets: Che et al. (2017); Gligic et al. (2020); Gupta et al. (2020); Amin-Nejad et al. (2020); Mahajan and Baths (2021); Guo et al. (2021); Zhu et al. (2021, 2022a); Ablimit et al. (2022a); Mirheidari et al. (2022); Hlédiková et al. (2022); Ilias and Askounis (2022b); El-Sappagh et al. (2022); Li et al. (2022b); Maclagan et al. (2023); Chen et al. (2023b); Farzana and Parde (2023); Cui et al. (2023); Chen et al. (2023a); Cai et al. (2023); Lin and Washington (2024) Fard et al. (2024).
@@ Appendix C Full Dataset Table
Table 2 (presented in a separate page due to margin requirements of the TACL format) showcases the dementia-related datasets encountered throughout our review. Datasets such as the Reagan Library and IMDB, though creatively used in some papers, are excluded as they are not dementia-specific. Private health records or unreleased clinical notes are excluded as well.
@@
| Farmington Heart Study | x | x | - | - | - | x | x | x | - | x | x |
- | - | x | - | - | - | x | - | - | - | - | - |